OpenStudy (anonymous):

How many solutions does this system of equations have? y + 5x = -12 -20x - 4y = -12 Question 8 options: One Infinite number None

jimthompson5910 (jim_thompson5910):

Notice that if you multiply equation by -4, you'll get -4(y+5x) = -4*(-12) -4y - 20x = 48 So we have the equations -20x-4y = 48 -20x-4y = -12 Add the equations to get -20x-4y = 48 -20x-4y = -12 -------------- 0x+0y = 36 0 = 36 ... which is a contradiction So there are no solutions

OpenStudy (anonymous):

o.k Thanks!!!

jimthompson5910 (jim_thompson5910):

You're welcome